OpenStudy (anonymous):

Please help?! I need step by step instruction when evaluating sin(Cos^-1(4/5)+Tan^1(3/4)). What is an identity used in the solution? a. cos2x = 1 - 2cos^2x b. cos2x = sin^2x - cos^2x c. sin2x = 1 - 2sin^2x d. sin2x = 2sin x cos x

OpenStudy (anonymous):

Do you know why one can say that \[ \alpha = \tan^{-1} (3/4)= \cos^{-1}(4/5) \]

OpenStudy (anonymous):

You have to compute \[ \sin(2 \alpha) \]

OpenStudy (anonymous):

\[ \sin(2\alpha) =2 \sin(\alpha) \cos(\alpha)= 2 \frac 3 5 \frac 4 5=\frac {24}{25} \] So the answer is d.

OpenStudy (anonymous):

Oh o.k. I get it now. Thank you for the explanation and help!

OpenStudy (anonymous):

yw